# Exponential mixing of random interval diffeomorphisms

### v1

@@ 2.3. Coupling time
Let us take $a>0$ sufficiently small to satisfy Proposition 1 and 3. Let $C$ be the constant given in Proposition 3. We insist also that the transition from $(0,a)$ to $(1-a,1)$ and from $(1-a,1)$ to $(0,a)$ is impossible in one step. By Proposition 1
@@
Since transition from $(0,a)$ to $(1-a,1)$ and from $(1-a,1)$ to $(0,a)$ is not possible in one step we have
@@
Let $A_{k}(z)\subseteq\Sigma$ be the set of $\omega$’s such that the moment of the first visit of $(f_{\omega}^{i}(z))$ in $[a,1-a]$ equals $k$. Let $B_{k}(z)$ be the set of $\omega$’s such that
@@ 2.4. Expansion after the first visit
The statement is obvious when $x,y\in[a,1-a]$ as the right-hand side of (2.12) is greater than 1, so we can consider only the situation when at least one of $x,y$ is outside $[a,1-a]$. Let us assume therefore that $x<y$ and $x<a$. To simplify the notation let us write $T(\omega)=T^{a}_{x,y}(\omega)$. Let $\delta$ be so small that $f^{n}_{\omega}(y)<ah^{-1}$ and $|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|\leq\frac{1}{2}$ for $n\leq K_{0}$ and $x,y$ such that $|x-y|<\delta$.
@@
for some $\zeta_{\omega}\in[x,y]$ and $\xi_{\omega}\in[0,x]$. By the choice of $\delta>0$ and the fact that $\omega\in\{T\leq K_{0}\}$, $T$ is the moment of the first visit of $f^{n}_{\omega}(x)$ in $[a,1-a]$, thus the value of $f^{T}_{\omega}(x)$ cannot be greater than $h^{-1}a$. Therefore $f^{T}_{\omega}(x)/x\leq h^{-1}a/x$ and
@@ 3.1. Proof of Theorem 1
By Lemma 4.1 in [GH17] $\Lambda<0$. Let $h$, $L^{\prime}$, $L^{\prime\prime}$ be the same as in the preceding section. Let $\alpha$, $C_{1}$, $\kappa$ be given in Proposition 1 and 4 respectively, and let $a>0$ be such that Propositions 1, 4, 5 are satisfied. Moreover, we assume
(3.1) $83\mu((0,a))\log\frac{C_{1}}{h}<\Lambda/8,\quad 3\mu((1-a,1))\log\frac{C_{1}}{h}<\Lambda/$ 8
@@
If $r^{-}(\omega,x)$, $r^{+}(\omega,x)$ are the numbers of visits of $(f^{n}_{\omega}(x))$ in $(0,a)$ and $(1-a,1)$, respectively, until $n\leq k_{0}$ then (3.2) implies
@@
We assume also $\beta$ to be so small that Proposition 5 is satisfied. By the mean value theorem and the continuity for every $x,y\in[a,1-a]$ with $x<y$ and $|x-y|$ sufficiently small (say less than $\delta$)
@@
In both cases $\tau$ is defined as the first moment greater or equal to either $k_{0}$ or $k_{1}$ that the trajectory visits $[a,1-a]$. Therefore Proposition 4 shows that $\sup_{x,y\in[a,1-a]}\mathbb{E}\exp(\kappa\tau_{x,y})<\infty$. Let
@@ 3.2. Proof of Corollary 1
Let $a>0$ be such that $\mathbb{E}|f^{n}_{\omega}(a)-f^{n}_{\omega}(1-a)|\leq Cq^{n}$, $n\geq 1$. Let $\varphi$ be a Hölder function with $|\varphi(x)-\varphi(y)|\leq L|x-y|^{\nu}$. We have by Theorem 1
@@
$=\bigg{|}U^{n}\varphi(x)-\int_{[0,1]}U^{n}\varphi(y)\mu_{\ast}(dy)\bigg{|}\leq L\int_{[0,1]}\mathbb{E}\big{|}f^{n}_{\omega}(x)-f^{n}_{\omega}(y)\big{|}^{\nu}\mu_{\ast}(dy)$
Let $\gamma>0$ be such that $\alpha\gamma-1/2<0$, where $\alpha$ is given in Proposition 1. Let $\xi_{n}=ac^{-\gamma n}$. Then $\mu_{\ast}((0,\xi_{n})\cup(1-\xi_{n},1))\leq 2M\xi_{n}^{\alpha}$, as $\mu_{\ast}\in\mathcal{P}_{M,\alpha}$ by Proposition 1. Let $R_{n}^{-}=\bigcap_{j=1}^{n/2}\{f^{j}_{\omega}(\xi_{n})<a\}$, $R_{n}^{+}=\bigcap_{j=1}^{n/2}\{f^{j}_{\omega}(\xi_{n})>1-a\}$. Using Proposition 1 again we get that $\mu_{\ast}(R_{n}^{-}\cup R_{n}^{+})\leq c^{n(1/2-\alpha\gamma)}$. Combining these two facts we reduce the proof to showing that
@@
[…]
@@
$\bigg{|}U^{n}\varphi(x)\bigg{|}\leq\sum_{l=1}^{\lfloor n/2\rfloor}\int_{C_{l}}\bigg{|}U^{n-l}\varphi(f^{l}_{\omega}(x))-U^{n-l}\varphi(a)\bigg{|}+\sum_{l=1}^{\lfloor n/2\rfloor}\mathbb{P}(C_{l})\bigg{|}U^{n-l}\varphi(a)\bigg{|}+\mathbb{P}(R_{n})$
Again by (3.6) and the Markov property the first term tends to zero like $C^{\nu}q^{\nu n/2}$. By (3.7) the second term decays not slower than $c^{n(1/2-\alpha\gamma)}+C^{\nu}q^{\nu n/2}+2M\xi_{n}^{\alpha}$. By Proposition 1 $\mathbb{P}(R_{n})\leq(\frac{a}{z})^{\alpha}c^{n/2}$, where $z=\min\{x,1-x\}$. This proves Corollary 1.
3.3. Proof of Corollary 2
We have
$\|U^{n}\varphi\|_{L^{p}(\mu_{\ast})}^{p}=\int_{[0,1]}|U^{n}\varphi(x)|^{p}\mu_{\ast}(dx)=\int_{[0,1]}\bigg{|}U^{n}\varphi(x)-\int_{[0,1]}\varphi(y)\mu_{\ast}(dy)\bigg{|}^{p}\mu_{\ast}(dx)$
$=\int_{[0,1]}\bigg{|}U^{n}\varphi(x)-\int_{[0,1]}\varphi(y)P^{n}\mu_{\ast}(dy)\bigg{|}^{p}\mu_{\ast}(dx)=\int_{[0,1]}\bigg{|}U^{n}\varphi(x)-\int_{[0,1]}U^{n}\varphi(y)\mu_{\ast}(dy)\bigg{|}^{p}\mu_{\ast}(dx)$
$=\int_{[0,1]}\bigg{|}\int_{[0,1]}\big{(}U^{n}\varphi(x)-U^{n}\varphi(y)\big{)}\mu_{\ast}(dy)\bigg{|}^{p}\mu_{\ast}(dx)\leq\int_{[0,1]}\int_{[0,1]}\big{|}U^{n}\varphi(x)-U^{n}\varphi(y)\big{|}^{p}\mu_{\ast}(dy)\mu_{\ast}(dx),$
where the last inequality is the Jensen inequality.
Let us define the sequence $\xi_{n}:=ac^{-\gamma n}$, where $\gamma>0$ is, as before, so small that $\alpha\gamma-1<0$. Let us decompose $[0,1]\times[0,1]$ as the sum of $K_{n}=[\xi_{n},1-\xi_{n}]\times[\xi_{n},1-\xi_{n}]$ and the remainder $R_{n}$. By Proposition 1 $\mu_{\ast}\in\mathcal{P}_{M,\alpha}$ and therefore $(\mu_{\ast}\otimes\mu_{\ast})(R_{n})\leq 4M\xi_{n}^{\alpha}$. Then
$\|U^{n}\varphi\|_{L^{p}(\mu_{\ast})}^{p}\leq\int_{[0,1]}\int_{[0,1]}\big{|}U^{n}\varphi(x)-U^{n}\varphi(y)\big{|}^{p}\mu_{\ast}(dy)\mu_{\ast}(dx)$
$\leq\int\int_{K_{n}}\bigg{(}L\mathbb{E}\big{|}f^{n}_{\omega}(x)-f^{n}_{\omega}(y)\big{|}^{\nu}\bigg{)}^{p}\mu_{\ast}(dy)\mu_{\ast}(dx)+\int\int_{R_{n}}\big{|}U^{n}\varphi(x)-U^{n}\varphi(y)\big{|}^{p}\mu_{\ast}(dy)\mu_{\ast}(dx)$
$\leq\bigg{(}L\mathbb{E}\big{|}f^{n}_{\omega}(\xi_{n})-f^{n}_{\omega}(1-\xi_{n})\big{|}^{\nu}\bigg{)}^{p}+\int\int_{R_{n}}2^{p}\|\varphi\|^{p}_{\infty}\mu_{\ast}(dy)\mu_{\ast}(dx)$
$\leq\bigg{(}L\mathbb{E}\big{|}f^{n}_{\omega}(\xi_{n})-f^{n}_{\omega}(1-\xi_{n})\big{|}^{\nu}\bigg{)}^{p}+2^{p+2}M\xi_{n}^{\alpha}\|\varphi\|^{p}_{\infty},$
where in the one but last inequality we used the fact that $\|U^{n}\varphi\|_{\infty}\leq\|\varphi\|_{\infty}$ and the triangle inequality.
Since $\xi_{n}^{\alpha}$ shrinks exponentially fast, we are left to estimate the first term in the last line. It easily follows by the same reasoning as in the end of the preceding proof that this term decays not slower than $c^{n(1/2-\alpha\gamma)}+C^{\nu}q^{\nu n/2}$. This completes the proof.

### v2

@@ 2.3. Coupling time
Let us take $a>0$ sufficiently small to satisfy Proposition 1 and 3. Let $C$ be the constant given in Proposition 3. We insist also that the transition from $(0,a)$ to (1a,1)(1-a,1) and from (1a,1)(1-a,1) to $(0,a)$ is impossible in one step. By Proposition 1
@@
Since transition from $(0,a)$ to (1a,1)(1-a,1) and from (1a,1)(1-a,1) to $(0,a)$ is not possible in one step we have
@@
Let $A_{k}(z)\subseteq\Sigma$ be the set of $\omega$’s such that the moment of the first visit of $(f_{\omega}^{i}(z))$ in [a,1a][a,1-a] equals $k$. Let $B_{k}(z)$ be the set of $\omega$’s such that
@@ 2.4. Expansion after the first visit
The statement is obvious when $x,y\in[a,1-a]$ as the right-hand side of (2.12) is greater than 1, so we can consider only the situation when at least one of $x,y$ is outside [a,1a][a,1-a]. Let us assume therefore that $x<y$ and $x<a$. To simplify the notation let us write $T(\omega)=T^{a}_{x,y}(\omega)$. Let $\delta$ be so small that $f^{n}_{\omega}(y)<ah^{-1}$ and $|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)|\leq\frac{1}{2}$ for $n\leq K_{0}$ and $x,y$ such that $|x-y|<\delta$.
@@
for some $\zeta_{\omega}\in[x,y]$ and $\xi_{\omega}\in[0,x]$. By the choice of $\delta>0$ and the fact that $\omega\in\{T\leq K_{0}\}$, $T$ is the moment of the first visit of $f^{n}_{\omega}(x)$ in [a,1a][a,1-a], thus the value of $f^{T}_{\omega}(x)$ cannot be greater than $h^{-1}a$. Therefore $f^{T}_{\omega}(x)/x\leq h^{-1}a/x$ and
@@ 3.1. Proof of Theorem 1
By Lemma 4.1 in [Gharaei_Homburg_2017] $\Lambda<0$. Let $h$, $^{\prime}, L′′L^{\prime\prime}$ be the same as in the preceding section. Let $\alpha$, $C_{1}$, $\kappa$ be given in Proposition 1 and 4 respectively, and let $a>0$ be such that Propositions 1, 4, 5 are satisfied. Moreover, we assume
(3.1) $3\mu((0,a))\log\frac{C_{1}}{h}<\Lambda/8,\quad 3\mu((1-a,1))\log\frac{C_{1}}{h}<\Lambda/8$
@@
If $r^{-}(\omega,x)$, $r^{+}(\omega,x)$ are the numbers of visits of $(f^{n}_{\omega}(x))$ in $(0,a)$ and (1a,1)(1-a,1), respectively, until $n\leq k_{0}$ then (3.2) implies
@@
We assume also $\beta$ to be so small that Proposition 5 is satisfied. By the mean value theorem and the continuity for every $x,y\in[a,1-a]$ with $x<y$ and |xy||x-y| sufficiently small (say less than $\delta$)
@@
In both cases $\tau$ is defined as the first moment greater or equal to either $k_{0}$ or $k_{1}$ that the trajectory visits [a,1a][a,1-a]. Therefore Proposition 4 shows that $\sup_{x,y\in[a,1-a]}\mathbb{E}\exp(\kappa\tau_{x,y})<\infty$. Let
@@ 3.2. Proof of Corollary 1
In the first step we are going to show that there exist $\tilde{C}\geq 1$, $\widetilde{q}\in(0,1)$ such that
(3.6) $\bigg|U^{n}\varphi(x)-\int_{0}^{1}\varphi\textrm{d}\mu_{\ast}\bigg|\leq\tilde{C}L\bigg(\frac{a}{\min\{x,1-x\}}\bigg)^{\alpha}\widetilde{q}^{\nu n}\quad\textrm{for $n\geq 1$.}$
for every Hölder function $\varphi$ with $|\varphi(x)-\varphi(y)|\leq L|x-y|^{\nu}$.
Let $a>0$ be such that $\mathbb{E}|f^{n}_{\omega}(a)-f^{n}_{\omega}(1-a)|\leq Cq^{n}$, $n\geq 1$. Let $\varphi$ be as above. We have by Theorem 1
@@
$\bigg|U^{n}\varphi(x)-\int_{[0,1]}U^{n}\varphi(y)\mu_{\ast}(dy)\bigg|\leq L\int_{[0,1]}\mathbb{E}\big|f^{n}_{\omega}(x)-f^{n}_{\omega}(y)\big|^{\nu}\mu_{\ast}(dy)$
Let $\gamma>0$ be such that $\alpha\gamma-1/2<0$, where $\alpha$ is given in Proposition 1. Let $\xi_{n}=ac^{-\gamma n}$. Then $\mu_{\ast}((0,\xi_{n})\cup(1-\xi_{n},1))\leq 2M\xi_{n}^{\alpha}$, as $\mu_{\ast}\in\mathcal{P}_{M,\alpha}$ by Proposition 1. Let $R_{n}^{-}=\bigcap_{j=1}^{n/2}\{\omega:f^{j}_{\omega}(\xi_{n})<a\}$, $R_{n}^{+}=\bigcap_{j=1}^{n/2}\{\omega:f^{j}_{\omega}(\xi_{n})>1-a\}$. Using Proposition 1 again we get that $\mathbb{P}(R_{n}^{-}\cup R_{n}^{+})\leq c^{n(1/2-\alpha\gamma)}$. Combining these two facts we reduce the proof to showing that
@@
For $x\not\in[a,1-a]$ it is convenient to assume without loss of generality that $\int_{0}^{1}\varphi d\mu_{\ast}=0$. Let
@@
$\bigg|U^{n}\varphi(x)\bigg|\leq\sum_{l=1}^{\lfloor n/2\rfloor}\int_{C_{l}}\bigg|U^{n-l}\varphi(f^{l}_{\omega}(x))\bigg|+\mathbb{P}(R_{n})\|\varphi\|_{\infty}$
By (3.8) the first term decays not slower than $c^{\frac{n}{2}(1/2-\alpha\gamma)}+C^{\nu}q^{\nu n/4}+2M\xi_{\lfloor n/2\rfloor}^{\alpha}$ (note (3.8) is uniform over $x\in[a,1-a]$). By Proposition 1 $\mathbb{P}(R_{n})\leq(\frac{a}{z})^{\alpha}c^{n/2}$, where $z=\min\{x,1-x\}$. This proves (3.6).
To pass to the assertion in the statement, fix two Lipschitz functions $\psi$, $\varphi$. Then (3.6) holds for $\varphi$ with $\nu=1$. Let $\zeta$ be such that $\zeta^{\alpha}\in(\widetilde{q},1)$. Let $K_{n}=[a\zeta^{n},1-a\zeta^{n}]\subseteq[0,1]$ and $R_{n}=[0,1]\setminus K_{n}$. We have by (3.6) and Proposition 1 again
$\bigg|\int_{0}^{1}U^{n}\varphi(x)\psi(x)d\mu_{\ast}(x)-\int_{0}^{1}\varphi(x)\textrm{d}\mu_{\ast}(x)\int_{0}^{1}\psi(x)\textrm{d}\mu_{\ast}(x)\bigg|\leq\int_{0}^{1}\bigg|U^{n}\varphi(x)-\int_{0}^{1}\varphi(x)\textrm{d}\mu_{\ast}(x)\bigg|\|\psi\|_{\infty}d\mu_{\ast}(x)$
$\int_{K_{n}}\bigg|U^{n}\varphi(x)-\int_{0}^{1}\varphi(x)\textrm{d}\mu_{\ast}(x)\bigg|\|\psi\|_{\infty}d\mu_{\ast}(x)+\int_{R_{n}}\bigg|U^{n}\varphi(x)-\int_{0}^{1}\varphi(x)\textrm{d}\mu_{\ast}(x)\bigg|\|\psi\|_{\infty}d\mu_{\ast}(x),$
$\leq\widetilde{C}\|\varphi\|_{\textrm{Lip}}\|\psi\|_{\textrm{Lip}}\bigg(\frac{a}{a\zeta^{n}}\bigg)^{\alpha}\widetilde{q}^{n}+4Ma^{\alpha}\zeta^{\alpha n}\|\varphi\|_{\textrm{Lip}}\|\psi\|_{\textrm{Lip}},$
which immediately implies the assertion by the choice of $\zeta$.